OpenStudy (anonymous):

4x + 5y = -9 4x – 8y = 4

ganeshie8 (ganeshie8):

subtract both equations vertically

ganeshie8 (ganeshie8):

4x + 5y = -9 4x – 8y = 4 ----------------------

ganeshie8 (ganeshie8):

wat do u get ?

OpenStudy (anonymous):

i really dont know how to do this im trying to get answers from people cause is grade what im doing on a website called ixl.com

OpenStudy (anonymous):

a daily grade

OpenStudy (anonymous):

PLZ HELP

ganeshie8 (ganeshie8):

hey brikalar giving out answers is a crime in openstudy. we discuss and learn stuff here, not just give answers oly :) so if you have time i can help u do these on ur own

ganeshie8 (ganeshie8):

subtract the equations vertically :- 4x + 5y = -9 4x – 8y = 4 ---------------------- 0 + 13x = -13 can u solve x from above ? :)